OpenStudy (anonymous):

A cheetah ran 704 feet in 8 seconds, and then another 328 feet in 4 seconds. What was the cheetah's average speed?

OpenStudy (luckycoins888):

Well, what is the total of feet the cheetah ran?

OpenStudy (anonymous):

do i do 704/8=88 then 328/4=82 then 82+88=170 then 170/2=85

OpenStudy (anonymous):

Take 104+ 328 then divid that by 2

OpenStudy (anonymous):

it cant be 380

OpenStudy (luckycoins888):

The way you did it is correct Makel.

OpenStudy (anonymous):

thxs

OpenStudy (luckycoins888):

Yep :)

OpenStudy (anonymous):

;)